OpenStudy (anonymous):

Use the rules of differentiation to find the derivative of the function y=.7x=3cosx

OpenStudy (anonymous):

check your statement.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

y=.7x=3cos x

OpenStudy (anonymous):

is it .7x=3cos x or .7x-3cosx

OpenStudy (anonymous):

it is y = .7x = 3cosx

OpenStudy (anonymous):

to me it is not correct. can you tell me the answer so that i can check if there is any change in statement?

OpenStudy (anonymous):

the answer isnt given

OpenStudy (anonymous):

can you try a different one and i can figure out the process?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=.7=-3 \sin x\]

OpenStudy (anonymous):

how would i do it for y= \[3\div \sqrt[4]{x}-7sinx\]

OpenStudy (anonymous):

\[y=\frac{ 3 }{ x ^{\frac{ 1 }{ 4 }} }-7 \sin x=3x ^{\frac{ -1 }{ 4 }}-7 \sin x\] \[\frac{ d }{ dx }\left( x^n \right)=nx ^{n-1}\] \[\frac{ d }{ dx }\sin x=\cos x\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=3*\left( -\frac{ 1 }{ 4 } \right)x ^{\frac{ -1 }{ 4 }-1}-7\cos x\] ?

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

yw